If the headlights on a car are separated by 1.3 m, how far down the road can they be resolved if the angular

resolution of the eye is about 5
Answer #1
We know that angular seperation ? = d/D
here d = seperation between the head lights = 1.3 m
? = 5* 10^-4 rad
D = DISTANCEBETWEEN THE OBSERVER AND SOURCE
THEN 5 * 10^-4 rad =1.3 m /D
SOLVE FOR D = 0.26 *104 m = 2600 m =2.6 km